Canal Inc. has an 1,000 par value bond with 15 years to maturity and a coupon rate of 14%, paid semiannually. The market rate on similar debt has now risen to 16%. What is the current price of this bond?
Answer #1

Rate = 16%/ 2 = 8%

Number of periods = 15 * 2 = 30

Coupon = ( 1,000 * 0.14) / 2 = 70

Current price = Coupon * [ 1 - 1 / ( 1 + r)n] / r + FV / ( 1 + r)n

Current price = 70 * [ 1 - 1 / ( 1 + 0.08)30] / 0.08 + 1000 / ( 1 + 0.08)30

Current price = 70 * 11.257783 + 99.37733

Current price = $887.42

Keys to use in a financial calculator:

FV 1000

N 30

PMT 70

I/Y 16

CPT PV